OpenStudy (anonymous):

Find the square. (5m + 7)^2 25m^2 + 35m - 49 25m^2 + 70m + 49 25m^2 + 70m - 49 25m^2 - 70m + 49 Won't let me pull this up on the calculator

OpenStudy (p0sitr0n):

the reason you cannot put this into your calculator is because m is a variable i.e. can take any value. you got to expand the square: (5m+7)^2 = 25m^2+70m+49

OpenStudy (p0sitr0n):

generally, \[(a+b)^2 = a^2+ 2ab+b^2\]

OpenStudy (anonymous):

answers 2 and 4 are the same answer what do I do if this happens

OpenStudy (p0sitr0n):

this means usually they are not the good ones xD

OpenStudy (anonymous):

there not the same lol